Brad purchases10 shares of AMZN on 60% initial margin for the price of $2100 per share. What will be his percent margin if the share price moves to $1800?
Answer #1

When the price is $2,100

His initial margin is 60%, which is 2,100 * 0.6 = 1,260

Borrowed amount = 2,100 - 1,260 = 840

When the price drops to $1,800, his margin = 1,800 - 840 = 960

His percent margin = 960/1,800 = 0.5333333333 = 53.33333333%

Can you please upvote? Thank You :-)